OpenStudy (shadowcain):

A recipe that makes 6 servings calls for 1/2 cup of butter. Nancy modifies the recipe so that it can serve 8. How many cups of butter does Nancy need?

OpenStudy (shadowcain):

1/2 ÷ 6 = 1/2 ÷ 6/1 = (1 × 1) / (6 × 2) = 1/12 per serving 1/12 x 8 = 1/12 × 8/1 = (1 × 8) / (12 × 1) = 8/12 = 2/3